# Conditional Wasserstein Distances with Applications in Bayesian OT Flow Matching

### v1

@@ 9 Conclusions
Inspired from applications in Bayesian inverse problems, we introduced conditional Wasserstein distances. We managed to rewrite these distances as expectations of the Wasserstein distances with respect to the observation. Therefore we are able to directly infer posterior guarantees when trained with the corresponding losses. Furthermore, we calculated the dual of the conditional Wasserstein-1 distance, when the probability measures are compactly supported and recovered well-known conditional Wasserstein GAN losses. We established corresponding velocity fields in the gradient flow theory and used our results to design a new Bayesian flow matching algorithm. Theoretically, improving Proposition 8 for non-discrete $P_{Y}$ distributions would be the next step. In particular one would need to show measurability of the "conditional Monge maps", which does not seem to be easy. Further, from a practical viewpoint, a clear use case would be in conditional domain translation, i.e., when the latent distribution is not a standard Gaussian, but given by some data distribution. There, finding a good OT matching and making use of our proposed framework could improve existing algorithms.
@@ Appendix A Counterexample for equality in equation 2
Figure 4: Visualization of the example. The blue dots belong to $P_{Y,X}$ and the red dots to $P_{Y,Z}$. The optimal coupling is the solid line, while the ”diagonal” coupling is the dotted one.
@@
where $\mathcal{S}_{n}$ is the set of permutations on $\{1,\ldots,n\}$, see [43, Proposition 2.1].
@@ Appendix B Proofs of Section 3
Proof of Proposition 1. i) First we show $\geq$. Let $\alpha_{y_{1},y_{2}}$ be the disintegration of some $\alpha\in\Gamma_{Y}^{4}(P_{Y,X},P_{Y,Z})$ with respect to $\pi^{1,3}_{\sharp}\alpha$. Then we obtain
$I(\alpha)$ $=\int_{(A\times B)^{2}}\|(y_{1},x_{1})-(y_{2},x_{2})\|^{p}\,\mathrm{d}\alpha(y_{1},x_{1},y_{2},x_{2})$ (33)
@@
Consequently, we have $I(\alpha)\geq\mathbb{E}_{y\sim P_{Y}}{\mathbb{E}}\big{[}W_{p}^{p}(P_{X|Y=y},P_{Z|Y=y})\big{]}$ and since $W_{p,Y}^{p}(P_{Y,X},P_{Y,Z})=\inf_{\alpha}I(\alpha)$, this gives assertion.
Now we prove the opposite direction $\leq$. For any $y\in A$, let $\alpha_{y}\in\Gamma(P_{X|Y=y},P_{Z|Y=y})$ be an optimal plan, i.e.
$\mathcal{W}_{p}^{p}(P_{X|Y=y},P_{Z|Y=y})=\int_{B^{2}}\|x_{1}-x_{2}\|^{p}\,\mathrm{d}\alpha_{y}(x_{1},x_{2}).$
This implies
$\int_{A}\mathcal{W}_{p}^{p}(P_{X|Y=y},P_{Z|Y=y})\,\mathrm{d}P_{Y}(y)=\int_{A}\int_{B^{2}}\|x_{1}-x_{2}\|^{p}\,\mathrm{d}\alpha_{y,y}(x_{1},x_{2})\mathrm{d}P_{Y}(y).$ (38)
@@
iii) Let $\alpha$ be defined by (11), i.e.,
By [6, Equation 5.3.1] the function $y\mapsto\int_{B^{2}}1_{\mathcal{S}}(y,x_{1},x_{2})\mathrm{d}\alpha_{y}$ is Borel measurable. Consequently also $y\mapsto\delta_{y}\times\alpha_{y}(\mathcal{S})$ is Borel measurable and thus $\alpha$ is well defined. Then (38) can be rewritten as
$\int_{A}\mathcal{W}_{p}^{p}(P_{X|Y=y},P_{Z|Y=y})\,\mathrm{d}P_{Y}(y)=\int_{(A\times B)^{2}}\|(y_{1},x_{1})-(y_{2},x_{2})\|^{p}\,\mathrm{d}\alpha(y_{1},x_{1},y_{2},x_{2}).$ (42)
for all Borel functions $f:A^{2}\to[0,+\infty]$, and the second one from
$=\int_{A\times B}f(y,x)\mathrm{d}P_{X|Y=y}(x)\,\mathrm{d}P_{Y}(y)$ (45)
$=\int_{A\times B}f(y,x)\,\mathrm{d}P_{Y,X}(y,x)$ (46)
for all Borel functions $f:A\times B\to[0,+\infty]$. The third equality follows analogously.
The final assertion follows from (42) and the equality relation (10). $\Box$
@@ Appendix C Proofs of Section 4
Proof of Proposition 4. Let $C_{b}=C_{b}(A\times B)$ be the space of continuous bounded functions on $A\times B$ and $\mathcal{S}$ the set of nonnegative finite Borel measures $\alpha$ on $(A\times B)^{2}$ which are supported at most on the y-diagonal. By [47, Section 1.2], we know that
@@
can become arbitrary large which contradicts $L(\alpha,-1,-1)\leq a$. Therefore the compactness of $A,B$ implies that $\mathcal{S}_{a}$ is a family of tight measures. By [11, Theorem 8.6.7], the set $\mathcal{S}_{a}$ is relatively compact in the weak topology. ∎
@@ Appendix D Proofs of Section 5
In summary, we see that $\alpha_{s,t}\in\Gamma_{2,Y}(\mu_{s},\mu_{t})$. Thus, we have
@@
Proof of Proposition 6. The statements i),ii) can given in [5, Lemma 17.3] (with $e\coloneqq e_{t}$, $\mu=\alpha$, $v=(y_{2},x_{2})-(y_{1},x_{1})$, $w=v_{t}$).
The continuity equation in iv) follow as in the proof of [5, Theorem 17.2].
@@
Thus, for any Borel measurable set $V\subseteq\mathbb{R}^{d}\times\mathbb{R}^{m}$ and $j\leq d$, we obtain by Part i) that
$\int_{V}(v_{t})_{j}\,\mathrm{d}\mu_{t}$ $=\int_{V}\,\mathrm{d}(e_{t})_{\sharp}((y_{2})_{j}-(y_{1})_{j})\alpha)=\int_{e_{t}^{-1}(V)}((y_{2})_{j}-(y_{1})_{j})\,\mathrm{d}\alpha=0.$ (75)
This implies $(v_{t}(y,x))_{j}=0$ for $\mu_{t}$-a.e. $(y,x)\in\mathbb{R}^{d}\times\mathbb{R}^{m}$. $\Box$
@@
    $\mu_{0},\mu_{1}$ are empirical measures with the same number of points and $T$ is an optimal map with associated optimal plan $\alpha\in\Gamma_{2}(\mu_{0},\mu_{1})$, or
@@
    $\mu_{0},\mu_{1}$ both admit densities $\rho_{0},\rho_{1}$ which are supported on open, convex, bounded, connected subsets $\Omega_{0},\Omega_{1}\subset\mathbb{R}^{m}$ on which they are bounded away from 00 and $\infty$. Assume further that $\rho_{0}\in C^{2}(\Omega_{0}),\rho_{1}\in C^{2}(\Omega_{1})$. Let $T$ be the optimal Monge map with associated optimal plan $\alpha\in\Gamma_{2}(\mu_{0},\mu_{1})$.
Let $\mu_{t}=(e_{t})_{\sharp}\alpha$ and $v_{t}\in L^{2}_{\mu_{t}}(\mathbb{R}^{m},\mathbb{R}^{m})$ with $v_{t}\mu_{t}=(e_{t})_{\sharp}(x_{2}-x_{1})\alpha$ which then satisfy the continuity equation. Then there exists a solution of the flow equation
@@
such that $\mu_{t}=\phi_{t,\sharp}\mu_{0}$. Furthermore, we have
@@
i): Let $\mu_{0}=\frac{1}{n}\sum_{i=1}^{n}\delta_{a_{i}},\mu_{1}=\frac{1}{n}\sum_{i=1}^{n}\delta_{b_{i}}$. The optimal plan is then $\alpha=\frac{1}{n}\sum_{i=1}^{n}\delta_{a_{i},T(a_{i})}$. Using $e_{t,\sharp}(x_{2}-x_{1})\alpha=v_{t}\mu_{t}$ and $\mu_{t}=\frac{1}{n}\sum_{i=1}^{n}\delta_{T_{t}(a_{i})}$ for $T_{t}(x)=(1-t)x+tT(x)$ we can conclude
@@
where $T_{t}=(1-t)\text{Id}+tT$. By Caffarelli’s regularity Theorem [51, Theorem 12.5, ii)], we get the existence of a unique Monge map $T\in C^{1}(\Omega_{1})$ mapping $\mu_{0}$ to $\mu_{1}$ and $U\in C^{1}(\Omega_{2})$ mapping $\mu_{1}$ to $\mu_{0}$. By [5, Theorem 5.2] we know that $T\circ U=\text{Id}$ on $\Omega_{2}$ and $U\circ T=\text{Id}$ on $\Omega_{2}$ and thus $T:\Omega_{1}\to\Omega_{2}$ is a $C^{1}$ diffeomorphism and in particular $\det(\nabla T)\neq 0$ on $\Omega_{1}$. Since we know by [6, Proposition 6.2.12] that $\nabla T$ is positive definite $\mu_{1}$ a.e. on $\Omega_{1}$ we can deduce from $\det(\nabla T)\neq 0$ that $\nabla T$ is positive definite on $\Omega_{1}$. Consequently for $T_{t}=(1-t)\text{Id}+tT$ we have that $\nabla T_{t}=(1-t)\text{Id}+t\nabla T$ is positive definite on $\Omega_{1}$ and thus the image of $\Omega_{1}$ under $T_{t}$ is open. Furthermore, we know by the proof of [6, Proposition 6.2.12] that $T_{t}$ as a Monge map from $\mu_{0}$ to $\mu_{t}$ is injective on all points where $\nabla T_{t}$ is positive definite, which is on the whole $\Omega_{1}$, and thus $T_{t}$ is a diffeomorphism onto its image. Consequently, it possesses a $C^{1}$ inverse $T_{t}^{-1}:T_{t}(\Omega_{1})\to\Omega_{1}$. Then $v_{t}:=(T-\text{Id})\circ T_{t}^{-1}:T_{t}(\Omega_{1})\to\mathbb{R}^{m}$ we have that $v_{t}$ is measurable since it is continuous on $T_{t}(\Omega_{1})$ and the same is true for $\phi_{t}=T_{t}:\Omega_{1}\to\mathbb{R}^{m}$. Furthermore, we have
@@
Since we can set $\phi_{t}(x)=x$ on $\mathbb{R}^{m}\setminus\Omega_{1}$ and $v_{t}(x)=0$ for $x\in\mathbb{R}^{m}\setminus T_{t}(\Omega_{1})$, we obtain the claim. ∎
@@
$\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}f(y,x)\,\mathrm{d}(e_{t})_{\sharp}((y_{2},x_{2})-(y_{1},x_{1})\alpha)$ $=\int_{(\mathbb{R}^{d}\times\mathbb{R}^{m})^{2}}f\circ e_{t}((y_{2},x_{2})-(y_{1},x_{1}))\,\mathrm{d}\alpha$ (80)
$=\frac{1}{n}\sum_{i=1}^{n}\int_{\mathbb{R}^{2m}}f\circ e_{t}((y_{i},x_{1}),(y_{i},x_{2}))\,\mathrm{d}\alpha_{y_{i}}$ (81)
@@
for $\mu_{0}$-a.e. $(y,x)\in\mathbb{R}^{d}\times\mathbb{R}^{m}$. Furthermore

### v2

@@ 9 Conclusions
Inspired from applications in Bayesian inverse problems, we introduced conditional Wasserstein distances. We managed to rewrite these distances as expectations of the Wasserstein distances with respect to the observation. Therefore we are able to directly infer posterior guarantees in expectation when trained with the corresponding losses. Furthermore, we calculated the dual of the conditional Wasserstein-1 distance, when the probability measures are compactly supported and recovered well-known conditional Wasserstein GAN losses. We established corresponding velocity fields in the gradient flow theory for geodesics and used our results to design a new Bayesian flow matching algorithm. Moreover we use an approximation for the conditional Wasserstein distance depending on a parameter $\beta$ and show numerically that, when using it for class conditional flow matching, the result does respect the classes for sufficiently large $\beta$. We achieve better FID than random Bayesian flow matching on Cifar10. Future work includes improving Proposition 8 for non-discrete $P_{Y}$ distributions, where the measurability of the "conditional Monge maps" is the main obstruction. Further, from a practical viewpoint, a clear use case would be in conditional domain translation, i.e., when the latent distribution is not a standard Gaussian, but given by some data distribution. There, finding a good OT matching and making use of our proposed framework could improve existing algorithms.
@@ Appendix A Counterexample for equality in equation 2
Figure 4: Visualization of the example. The blue dots belong to $P_{Y,X}$ and the red dots to $P_{Y,Z}$. The optimal coupling is the solid line, while the optimal conditional coupling is the dotted one.
@@
where $\mathcal{S}_{n}$ is the set of permutations on $\{1,\ldots,n\}$, see [48, Proposition 2.1].
@@ Appendix B Proofs of Section 3
Proof of Proposition 1. iii) Let $\alpha$ be defined by (11) which was already used in the proof of [35, Theorem 2] , i.e.,
By [6, Equation 5.3.1] the function $y\mapsto\int_{B^{2}}1_{\mathcal{S}}(y,x_{1},x_{2})\mathrm{d}\alpha_{y}$ is Borel measurable. Consequently also $y\mapsto\delta_{y}\times\alpha_{y}(\mathcal{S})$ is Borel measurable and thus $\alpha$ is well defined.
for all Borel functions $f:A^{2}\to[0,+\infty]$, and the second one from
$=\int_{A\times B}f(y,x)\mathrm{d}P_{X|Y=y}(x)\,\mathrm{d}P_{Y}(y)$ (38)
$=\int_{A\times B}f(y,x)\,\mathrm{d}P_{Y,X}(y,x)$ (39)
for all Borel functions $f:A\times B\to[0,+\infty]$. The third equality follows analogously. The optimality of $\alpha$ for $W_{p,Y}(P_{Y,X},P_{Y,Z})$ follows from (47) which we show below.
i) First we show $\geq$. Let $\alpha_{y_{1},y_{2}}$ be the disintegration of some $\alpha\in\Gamma_{Y}^{4}(P_{Y,X},P_{Y,Z})$ with respect to $\pi^{1,3}_{\sharp}\alpha$. Then we obtain
$I(\alpha)$ $:=\int_{(A\times B)^{2}}\|(y_{1},x_{1})-(y_{2},x_{2})\|^{p}\,\mathrm{d}\alpha(y_{1},x_{1},y_{2},x_{2})$ (40)
@@
Consequently, we have $I(\alpha)\geq\mathbb{E}_{y\sim P_{Y}}\big{[}W_{p}^{p}(P_{X|Y=y},P_{Z|Y=y})\big{]}$ and since $W_{p,Y}^{p}(P_{Y,X},P_{Y,Z})=\inf_{\alpha}I(\alpha)$, this gives the assertion.
Now we prove the opposite direction $\leq$. Let $\alpha\coloneqq\int_{A}\mathrm{d}\delta_{y_{1}}(y_{2})\,\mathrm{d}\alpha_{y_{1}}(x_{1},x_{2})\mathrm{d}P_{Y}(y_{1})$ be as in $iii)$ i.e. $W_{p}^{p}(P_{X|Y=y},P_{Z|Y=y})=\int_{B^{2}}\|x_{1}-x_{2}\|^{p}\mathrm{d}\alpha_{y}$. Then
$I(\alpha)$ $=\int_{(A\times B)^{2}}\|(y_{1},x_{1})-(y_{2},x_{2})\|^{p}\mathrm{d}\delta_{y_{1}}(y_{1})\mathrm{d}\alpha_{y_{1}}(x_{1},x_{2})$ (45)
$=\int_{A}\int_{B^{2}}\|x_{1}-x_{2}\|^{p}\mathrm{d}\alpha_{y}\mathrm{d}P_{Y}=\mathbb{E}_{y\sim P_{Y}}\big{[}W_{p}^{p}(P_{X|Y=y},P_{Z|Y=y})\big{]}$ (46)
which gives the assertion and we conclude $\mathbb{E}_{y\sim P_{Y}}\big{[}W_{p}^{p}(P_{X|Y=y},P_{Z|Y=y})\big{]}=W_{p,Y}(P_{Y,X},P_{Y,Z})$. We also obtain
$I(\alpha)\leq\mathbb{E}_{y\sim P_{Y}}\big{[}W_{p}^{p}(P_{X|Y=y},P_{Z|Y=y})\big{]}=W_{p,Y}(P_{Y,X},P_{Y,Z})$ (47)
which shows that the coupling $\alpha$ from $iii)$ is optimal for $W_{p,Y}(P_{Y,X},P_{Y,Z})$.
@@
$\Box$
@@ Appendix C Proofs of Section 4
Proof of Proposition 4. Let $C_{b}=C_{b}(A\times B)$ be the space of continuous bounded functions on $A\times B$ and $\mathcal{S}$ the set of nonnegative finite Borel measures $\alpha$ on $(A\times B)^{2}$ which are supported at most on the y-diagonal. By [52, Section 1.2], we know that
@@
can become arbitrary large which contradicts $L(\alpha,-1,-1)\leq a$. Therefore the compactness of $A,B$ implies that $\mathcal{S}_{a}$ is a family of tight measures. By [12, Theorem 8.6.7], the set $\mathcal{S}_{a}$ is relatively compact in the weak topology. ∎
@@ Appendix D Proofs of Section 5
In summary, we see that $\alpha_{s,t}\in\Gamma_{Y}^{4}(\mu_{s},\mu_{t})$. Thus, we have
@@
Proof of Proposition 6. The proof of $i),ii),iv)$ is almost identical to the proofs of [5, Theorem 17.2, Lemma 17.3]. For the convenience of the reader and because we also need measurability of $v_{t}$ in $t$ we include a slight adaptation of their proof. We let $e:[0,1]\times\mathbb{R}^{d}\times\mathbb{R}^{m}\times\mathbb{R}^{d}\times\mathbb{R}^{m}\to[0,1]\times\mathbb{R}^{d}\times\mathbb{R}^{m}$ be defined by $e(t,y_{1},x_{1},y_{2},x_{2})=(t,(1-t)(y_{1},x_{1})+t(y_{2},x_{2}))$. For $\mathcal{L}$ the Lebesgue measure we have that
$\int_{[0,1]\times\mathbb{R}^{d+m}}f(t,y,x)\mathrm{d}e_{\sharp}\left(\mathcal{L}\otimes\alpha\right)$ $=\int_{[0,1]}\int_{\mathbb{R}^{2d+2m}}f(t,e_{t}(y_{1},x_{1},y_{2},x_{2}))\mathrm{d}\alpha\mathrm{d}t$ (75)
$=\int_{[0,1]}\int_{\mathbb{R}^{d+m}}f(t,y,x)\mathrm{d}e_{t,\sharp}\alpha\mathrm{d}t=\int_{[0,1]}\int_{\mathbb{R}^{d+m}}f(t,y,x)\mathrm{d}\mu_{t}\mathrm{d}t$ (76)
for every bounded $\mathcal{L}\otimes\alpha$ measurable function $f$. In particular using $f=1_{B}$ for a Borel set $B\in\mathcal{B}(\mathbb{R}^{d}\times\mathbb{R}^{m})$ we see that $\mu_{t}:[0,1]\times\mathcal{B}\to\mathbb{R}$ is a Markov kernel, i.e. the measure $\int_{0}^{1}\mu_{t}\mathrm{d}t$ is well defined, and $e_{\sharp}\left(\mathcal{L}\otimes\alpha\right)=\int_{0}^{1}\mu_{t}\mathrm{d}t$. Hence [5, Lemma 17.3] (with $e\coloneqq e$, $\mu=\mathcal{L}\otimes\alpha$, $v=(y_{2},x_{2})-(y_{1},x_{1})$, $w=v$) implies that there exists $v\in L^{2}_{\int\mu_{t}\mathrm{d}t}([0,1]\times\mathbb{R}^{d}\times\mathbb{R}^{m})$ such that
$e_{\sharp}\left(((y_{2},x_{2})-(y_{1},x_{1}))\mathcal{L}\otimes\alpha\right)=v\int_{0}^{1}\mu_{t}\mathrm{d}_{t}$ (77)
and we can choose a Borel measurable representative of $v.$ Since for test functions $f$ we have the following identities
$\int_{[0,1]\times\mathbb{R}^{d}\times\mathbb{R}^{m}}f\mathrm{d}e_{\sharp}\left(((y_{2},x_{2})-(y_{1},x_{1}))\mathcal{L}\otimes\alpha\right)$ $=\int_{0}^{1}\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}f\mathrm{d}e_{t,\sharp}\left(((y_{2},x_{2})-(y_{1},x_{1}))\alpha\right)\mathrm{d}t$ (78)
$\int_{[0,1]\times\mathbb{R}^{d}\times\mathbb{R}^{m}}f\mathrm{d}\left(v\int_{0}^{1}\mu_{t}\mathrm{d}_{t}\right)$ $=\int_{0}^{1}\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}fv_{t}\mathrm{d}\mu_{t}\mathrm{d}t$ (79)
we obtain from (77) that $e_{t,\sharp}\left(((y_{2},x_{2})-(y_{1},x_{1}))\alpha\right)=v(t,y,x)\mu_{t}$ for $\mathcal{L}$ a.e. $t\in[0,1]$ which shows $i)$. By [5, Lemma 17.3] we obtain that
$\|v_{t}(y,x)\|_{L^{2}_{\mu_{t}}}\leq\|((y_{2},x_{2})-(y_{1},x_{1}))\|_{L^{2}_{\alpha}}=W_{2,Y}(\mu_{0},\mu_{1})$
for a.e. $t\in[0,1]$ and hence we can conclude $ii)$.
@@
Thus, for any Borel measurable set $V\subseteq[0,1]\times\mathbb{R}^{d}\times\mathbb{R}^{m}$ and $j\leq d$, we obtain by Part i) that
$\int_{V}v_{j}\,\mathrm{d}\mu_{t}\mathrm{d}t$ $=\int_{V}\,\mathrm{d}e_{\sharp}((y_{2})_{j}-(y_{1})_{j})(\mathcal{L}\otimes\alpha))=\int_{0}^{1}\int_{\tilde{V}}((y_{2})_{j}-(y_{1})_{j})\,\mathrm{d}\alpha\mathrm{d}t=0$ (83)
where $\tilde{V}=(\pi^{t})^{-1}(e^{-1}(V))$. This implies $(v(t,y,x))_{j}=0$ for $\int_{0}^{1}\mu_{t}\mathrm{d}t$-a.e. $(t,y,x)\in[0,1]\times\mathbb{R}^{d}\times\mathbb{R}^{m}$ and we can choose a Borel measurable representative of $v$ such that $v_{j}=0$ for all $(t,y,x)\in[0,1]\times\mathbb{R}^{d}\times\mathbb{R}^{m}$
Next we proof $iv)$. Let $\varphi\in C_{c}^{\infty}((0,1)\times\mathbb{R}^{d}\times\mathbb{R}^{m})$. Then by the chain rule
$\frac{\partial}{\partial t}(\varphi(t,e_{t}))=\left(\frac{\partial}{\partial t}\varphi\right)\circ(t,e_{t})+\langle\nabla_{y,x}\varphi(t,e_{t}),(y_{2},x_{2})-(y_{1},x_{1})\rangle.$
Consequently
$\int_{0}^{1}\int_{\mathbb{R}^{d+m}}\frac{\partial}{\partial t}\varphi\mathrm{d}\mu_{t}\mathrm{d}t$ $=\int_{0}^{1}\int_{\mathbb{R}^{2d+2m}}\left(\frac{\partial}{\partial t}\varphi\right)\circ(t,e_{t})\mathrm{d}\alpha\mathrm{d}t$ (84)
$=\int_{0}^{1}\int_{\mathbb{R}^{2d+2m}}\frac{\partial}{\partial t}(\varphi(t,e_{t}))-\langle\nabla_{y,x}\varphi(t,e_{t}),(y_{2},x_{2})-(y_{1},x_{1})\rangle\mathrm{d}\alpha\mathrm{d}t$ (85)
$=\int_{0}^{1}\frac{\partial}{\partial t}\int_{\mathbb{R}^{2d+2m}}\varphi(t,e_{t}))\mathrm{d}\alpha\mathrm{d}t$ (86)
$-\int_{0}^{1}\int_{\mathbb{R}^{2d+2m}}\langle\nabla_{y,x}\varphi(t,e_{t}),(y_{2},x_{2})-(y_{1},x_{1})\rangle\mathrm{d}\alpha\mathrm{d}t$ (87)
$=0-\int_{0}^{1}\int_{\mathbb{R}^{d+m}}\langle\nabla_{y,x}\varphi,e_{t,\sharp}(y_{2},x_{2})-(y_{1},x_{1})\mathrm{d}\alpha\rangle\mathrm{d}t$ (88)
$=-\int_{0}^{1}\int_{\mathbb{R}^{d+m}}\langle\nabla_{y,x}\varphi,v_{t}\rangle\mathrm{d}\mu_{t}\mathrm{d}t$ (89)
where we used
$\int_{0}^{1}\frac{\partial}{\partial t}\int_{\mathbb{R}^{2d+2m}}\varphi(t,e_{t}))\mathrm{d}\alpha\mathrm{d}t=\int_{\mathbb{R}^{2d+2m}}\varphi(1,e_{1}))\mathrm{d}\alpha-\int_{\mathbb{R}^{2d+2m}}\varphi(0,e_{0}))\mathrm{d}\alpha=0$
since $\varphi(1,\cdot)=\varphi(0,\cdot)=0$ because $\varphi$ is compactly supported on $(0,1)\times\mathbb{R}^{d}\times\mathbb{R}^{m}$. $\Box$
@@
    $\mu_{0},\mu_{1}$ are empirical measures with the same number of points and $T$ is an optimal map with associated optimal plan $\alpha\in\Gamma(\mu_{0},\mu_{1})$, or
@@
    $\mu_{0},\mu_{1}$ both admit densities $\rho_{0},\rho_{1}$ which are supported on open, convex, bounded, connected subsets $\Omega_{0},\Omega_{1}\subset\mathbb{R}^{m}$ on which they are bounded away from 00 and $\infty$. Assume further that $\rho_{0}\in C^{2}(\Omega_{0}),\rho_{1}\in C^{2}(\Omega_{1})$. Let $T$ be the optimal Monge map with associated optimal plan $\alpha\in\Gamma(\mu_{0},\mu_{1})$.
Let $\mu_{t}=(e_{t})_{\sharp}\alpha$ and $v_{t}\in L^{2}_{\mu_{t}}(\mathbb{R}^{m},\mathbb{R}^{m})$ with $v_{t}\mu_{t}=(e_{t})_{\sharp}(x_{2}-x_{1})\alpha$ which then satisfy the continuity equation. Then there is a Borel measurable representative $v_{t}$ such that there exists a solution of the flow equation
@@
and $\mu_{t}=\phi_{t,\sharp}\mu_{0}$. Furthermore, we have
@@
i): Let $\mu_{0}=\frac{1}{n}\sum_{i=1}^{n}\delta_{a_{i}},\mu_{1}=\frac{1}{n}\sum_{i=1}^{n}\delta_{b_{i}}$ and let $T$ be a optimal map. The associated optimal plan is then $\alpha=\frac{1}{n}\sum_{i=1}^{n}\delta_{a_{i},T(a_{i})}$. Using $e_{t,\sharp}(x_{2}-x_{1})\alpha=v_{t}\mu_{t}$ and $\mu_{t}=\frac{1}{n}\sum_{i=1}^{n}\delta_{T_{t}(a_{i})}$ for $T_{t}(x)=(1-t)x+tT(x)$ we can conclude
@@
where $T_{t}=(1-t)\text{Id}+tT$. By Caffarelli’s regularity Theorem [56, Theorem 12.5, ii)], we get the existence of a unique Monge map $T\in C^{1}(\Omega_{0})$ mapping $\mu_{0}$ to $\mu_{1}$ and $U\in C^{1}(\Omega_{1})$ mapping $\mu_{1}$ to $\mu_{0}$. By [5, Theorem 5.2] we know that $T\circ U=\text{Id}$ on $\Omega_{1}$ and $U\circ T=\text{Id}$ on $\Omega_{0}$ and thus $T:\Omega_{0}\to\Omega_{1}$ is a $C^{1}$ diffeomorphism and in particular $\det(\nabla T)\neq 0$ on $\Omega_{0}$. Since we know by [6, Proposition 6.2.12] that $\nabla T$ is positive definite $\mu_{1}$ a.e. on $\Omega_{0}$ we can deduce from $\det(\nabla T)\neq 0$ that $\nabla T$ is positive definite on $\Omega_{0}$. Consequently for $T_{t}=(1-t)\text{Id}+tT$ we have that $\nabla T_{t}=(1-t)\text{Id}+t\nabla T$ is positive definite on $\Omega_{0}$ and thus the image of $\Omega_{0}$ under $T_{t}$ is open. Furthermore, we know by the proof of [6, Proposition 6.2.12] that $T_{t}$ as a Monge map from $\mu_{0}$ to $\mu_{t}$ is injective on all points where $\nabla T_{t}$ is positive definite, which is on the whole $\Omega_{0}$, and thus $T_{t}$ is a diffeomorphism onto its image. Consequently, it possesses a $C^{1}$ inverse $T_{t}^{-1}:T_{t}(\Omega_{0})\to\Omega_{0}$. Furthermore $(t,x)\mapsto(t,T_{t}(x))$ is an bijective Borel map from $[0,1]\times\Omega_{0}$ onto its Borel measurable image which we denote by $\Omega\subset[0,1]\times\mathbb{R}^{m}$. Thus $T_{t}^{-1}$ is a Borel measurable map from $\Omega\to\Omega_{0}$. Then for $v_{t}:=(T-\text{Id})\circ T_{t}^{-1}:T_{t}(\Omega_{0})\to\mathbb{R}^{m}$ we have that $v_{t}$ is Borel measurable on $\Omega$ and also $\phi_{t}=T_{t}:\Omega_{0}\to\mathbb{R}^{m}$ is Borel measurable. Furthermore, we have
@@
Since we can set $\phi_{t}(x)=x$ on $\mathbb{R}^{m}\setminus\Omega_{0}$ and $v_{t}(x)=0$ for $x\in\mathbb{R}^{m}\setminus T_{t}(\Omega_{0})$, we obtain the claim. ∎
@@
$\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}f(y,x)\,\mathrm{d}(e_{t})_{\sharp}((y_{2},x_{2})-(y_{1},x_{1})\alpha)$ $=\int_{(\mathbb{R}^{d}\times\mathbb{R}^{m})^{2}}f\circ e_{t}\cdot((y_{2},x_{2})-(y_{1},x_{1}))\,\mathrm{d}\alpha$ (94)
$=\frac{1}{n}\sum_{i=1}^{n}\int_{\mathbb{R}^{2m}}f\circ e_{t}\cdot((y_{i},x_{1}),(y_{i},x_{2}))\,\mathrm{d}\alpha_{y_{i}}$ (95)
@@
$=\frac{1}{n}\sum_{i=1}^{n}\int_{\mathbb{R}^{2m}}f\mathrm{d}(\tilde{e}_{t})_{\sharp}(0,x_{2}-x_{1})\alpha_{y_{i}}$ (97)
@@
for $\mu_{0}$-a.e. $(y,x)\in\mathbb{R}^{d}\times\mathbb{R}^{m}$. Note that since $\tilde{v}_{t,y_{i}}$ is Borel measurable on $[0,1]\times\mathbb{R}^{m}$ we can assume that $\tilde{v}_{t}$ is Borel measurable on $[0,1]\times\mathbb{R}^{d}\times\mathbb{R}^{m}$ and similarly for $\phi_{t}$. Furthermore